# Stochastic SIR model with individual heterogeneity and infection-age dependent infectivity on large non-homogeneous random graphs

### v1

@@ A.2. Almost everywhere continuity
The measurability of the set $\Phi[\delta,\eta]$ can be verified through the following definition, which happens to be equivalent to the one given in (A.2):
$\Phi[\delta,\eta]=\cup_{\{q\in{\mathbf{Q}}_{+},m\geq 1\}}\bar{\omega}^{-1}([0,q])^{\eta}\cap\bar{\omega}^{-1}([q+\delta+2^{-m},\omega^{*}])^{\eta},$
so that $A^{\eta}$ denotes the $\eta$-vicinity of $A$.
The measurability of the sets $\Psi[\delta,\eta]$ is a bit more technical than the one of $\Phi[\delta,\eta]$, yet follows similar ideas. We exploit here the density in $L^{1}(\bar{\mu}_{X})$ of a countable subset $\mathfrak{Q}$, according to [16, Proposition 3.4.5]. Let $\varepsilon_{m}=2^{-m}$ and $\mathcal{W}:x\in\mathbf{\mathbb{X}}\mapsto\omega(x,.)$. Then, we claim that the following definition is equivalent to the one of $\Psi[\delta,\eta]$ in (A.1):
$\overline{\Psi}[\delta,\eta]=\cup_{\{m\geq 1\}}\cup_{\{(\omega_{1},\omega_{2})\in\mathfrak{Q}^{2}\cap\mathfrak{D}_{\delta,m}\}}\mathcal{W}^{-1}\Big{(}B_{L^{1}(\bar{\mu}_{X})}(\omega_{1},\varepsilon_{m})\Big{)}^{\eta}\cap\mathcal{W}^{-1}\Big{(}B_{L^{1}(\bar{\mu}_{X})}(\omega_{2},\varepsilon_{m})\Big{)}^{\eta}\,,$
where the couple $(\omega_{1},\omega_{2})\in\mathfrak{Q}^{2}$ belong to $\mathfrak{D}_{\delta,m}$ provided that
$\int_{\mathbf{\mathbb{X}}}|\omega_{1}-\omega_{2}|\,\bar{\mu}_{X}\geq\delta+2\varepsilon_{m}.$ (A.6)
It is clear from this definition that $\overline{\Psi}[\delta,\eta]$ is measurable. We next prove the equality with $\Psi[\delta,\eta]$. If $x\in\overline{\Psi}[\delta,\eta]$, then it means that there exists $m\geq 1$, $(\omega_{1},\omega_{2})\in\mathfrak{Q}^{2}\cap\mathfrak{D}_{\delta,m}$ and $(x_{1},x_{2})\in B(x,\eta)^{2}$ such that
$\int_{\mathbf{\mathbb{X}}}|\omega(x_{1},.)-\omega_{1}|\,\bar{\mu}_{X}<\varepsilon_{m},\quad\int_{\mathbf{\mathbb{X}}}|\omega(x_{2},.)-\omega_{2}|\,\bar{\mu}_{X}<\varepsilon_{m}.$
This entails the following property from which $x\in\Psi[\delta,\eta]$ is a direct consequence:
$\int_{\mathbf{\mathbb{X}}}|\omega(x_{1},.)-\omega(x_{2},.)|\,\bar{\mu}_{X}>\delta.$ (A.7)
Reciprocally, if we assume that there exist a couple $(x_{1},x_{2})\in B(x,\eta)^{2}$ which satisfies (A.7), then we can first choose $m\geq 1$ such that
$\int_{\mathbf{\mathbb{X}}}|\omega(x_{1},.)-\omega(x_{2},.)|\,\bar{\mu}_{X}\geq\delta+4\varepsilon_{m}.$
We can next identify two elements $(\omega_{1},\omega_{2})\in\mathfrak{Q}^{2}$ in respectively $B_{L^{1}(\bar{\mu}_{X})}(\omega(x_{1},.),\varepsilon_{m})$ and $B_{L^{1}(\bar{\mu}_{X})}(\omega(x_{2},.),\varepsilon_{m})$. By construction, $(\omega_{1},\omega_{2})\in\mathfrak{D}_{\delta,m}$, $x_{1}\in\mathcal{W}^{-1}\Big{(}B_{L^{1}(\bar{\mu}_{X})}(\omega_{1},\varepsilon_{m})\Big{)}$ and similarly for $x_{2}$ (with $\omega_{2}$ instead of $\omega_{1}$). This entails that $x\in\Psi[\delta,\eta]$. From this we have concluded that $\Psi[\delta,\eta]=\overline{\Psi}[\delta,\eta]$, so that the former is indeed measurable. ∎
@@
In the degenerate case where $\nu\equiv 0$, the convergence of $\langle\mu^{N},\varepsilon^{N}\rangle$ to 0 can be directly deduced with the upper-bound of $\varepsilon^{N}$ by $\|k\|_{\infty}\cdot\nu^{N}(\mathcal{Y})$ which converges to 0. In the degenerate case where $\mu\equiv 0$, it suffices to take $3\|k\|_{\infty}\cdot\nu(\mathcal{Y})$ as the uniform upper-bound of $\varepsilon^{N}$ for any $N$ sufficiently large, as $\mu^{N}(\mathcal{X})$ then tends to 0. In the following, we can thus assume that both $\nu(\mathcal{Y})>0$ and $\mu(\mathcal{X})>0$.
The irregularities of $k$ will be located through the following subset $\mathcal{G}[\delta,\eta]$ of $\mathcal{X}\times\mathcal{Y}$, defined for any $\delta,\eta>0$ similarly as in (A.2):
@@
while $B[(x,y);\eta]$ denotes the open ball centered in $(x,y)$ of radius $\eta$. The size $\eta$ of the vicinities in (A.9) shall be considered sufficiently small to ensure that discrepancies of order $\delta$ are exceptional. The measurability of $\mathcal{G}[\delta,\eta]$, for any $\delta,\eta>0$, is proved as in Lemma A.3.
@@
Since $k$ is continuous $\mu\otimes\nu$ almost everywhere, in particular, $k(x,.)$ is continuous $\nu$ almost everywhere for $x$ on a measurable set $\mathcal{A}\subset\mathcal{X}$ such that $\mu(\mathcal{A})=1$. For any $x\in\mathcal{A}$, thanks to the Portmanteau theorem, see e.g. [31], Subsection IV.3a on the ”Weak Convergence of Probability Measures”, $\varepsilon^{N}(x)$ converges to 0. Remark as compared to the classical version of Portmanteau theorem that we allow $\nu^{N}$ and $\nu$ to be general non-negative finite measure rather than probability measures, given that the proof is not difficult to adapt for this setting. As a consequence of Lebesgue’s dominated convergence theorem, recalling that $\varepsilon^{N}$ is bounded (by 1 under our assumption), we deduce
@@
Step 2: Convergence of $[\mu\otimes\nu](\mathcal{G}[\delta,\eta])$ and $[\mu\otimes\nu^{N}](\mathcal{G}[\delta,\eta])$ to zero
@@
Note also that the sets $\mathcal{G}[\delta,\eta]$ are increasing as $\delta$ decreases and non-increasing as $\eta$ decreases. Therefore, due to the fact that $k$ is $\mu\otimes\nu$ almost everywhere continuous, the following convergence to zero holds for any $\delta$:
@@
$[\mu\otimes\nu^{N}](\mathcal{G}[\delta,\eta])\leq[\mu\otimes\nu](\mathcal{G}[\delta,\eta]^{\eta})+\eta,$ (A.13)
@@
$\begin{split}\Big{|}\int_{\mathcal{G}_{x}[\delta,\eta]^{c}}|k(x,y)-k(x^{\prime},y)|\,[\nu^{N}-\nu](dy)\Big{|}&\leq[\nu^{N}\vee\nu](\mathcal{G}_{x}[\delta,\eta]^{c})\cdot\delta\\ &\leq 2\delta\cdot\nu(\mathcal{Y})\,,\end{split}$ (A.17)
@@
$[\nu\vee\nu^{N}](\mathcal{G}_{x}[\delta,\eta])\geq\delta.$
@@
Step 4: Proof of Proposition 5.9 in the particular case where $(\mu^{N})$ and $(\nu^{N})$ are deterministic.
@@
We choose $\eta\in(0,\delta)$ sufficiently small thanks to Step 2, to ensure both that $[\mu\otimes\nu](\mathcal{G}[\delta,\eta])$ is smaller than $\delta^{2}/2$ and similarly for $[\mu\otimes\nu^{N}](\mathcal{G}[\delta,\eta])$ for any $N$ sufficiently large. Thanks to the Markov inequality, we have
@@
It remains to combine this result with (A.19) and (A.22) to conclude the proof that $(\langle\mu^{N},\varepsilon^{N}\rangle)$ tends to 0 as $N$ tends to infinity, since $\delta$ can be taken arbitrarily small. This concludes the proof of Proposition 5.9 in the particular case where the sequences $(\mu^{N})$ and $(\nu^{N})$ are deterministic.
@@
For this final step, we no longer require the sequences $(\mu^{N})$ and $(\nu^{N})$ to be a priori deterministic, though our approach consists in referring to this convenient situation. We exploit [35, Lemma 4.2] to relate the convergence of probability to a.s. convergence of sequence extractions. Let $(N[k])_{k\geq 1}\in\mathbb{N}^{\mathbb{N}}$ be an increasing sequence. Since $\mu^{N}$ and $\nu^{N}$ converge in probability, we can extract a subsequence $(\check{N}[\ell])_{\ell\geq 1}=(N[K[\ell]])_{\ell\geq 1}$ from this sequence such that $(\check{\mu}^{\ell})=(\mu^{\check{N}[\ell]})$ and $(\check{\nu}^{\ell})=(\nu^{\check{N}[\ell]})$ converge a.s. respectively to $\mu$ and $\nu$. On this event of probability 1, we deduce from Step 4 that $\langle\check{\mu}^{\ell},\check{\varepsilon}^{\ell}\rangle$ converges to 0 as $\ell$ tends to infinity. Since the convergence of such an extraction of the sequence $(\langle\mu^{N},\varepsilon^{N}\rangle)$ holds whatever the initial extraction, this concludes the proof of Proposition 5.9 in that $(\langle\mu^{N},\varepsilon^{N}\rangle)$ converges in probability to 0. ∎

### v2

@@ A.2. Almost everywhere continuity
For any $y_{1},y_{2}\in\mathcal{Y}$:
$\begin{split}&|\int_{0}^{t}k(y_{1},a)\,F(y_{1},a)\,\mathrm{d}a-\int_{0}^{t}k(y_{2},a)\,F(y_{2},a)\,\mathrm{d}a|\\ &\quad\leq\|F\|_{\infty}\,\|k(y_{1},.)-k(y_{2},.)\|_{L^{1}(Leb)}+t\,\|k\|_{\infty}\,\|F(y_{1},.)-F(y_{2},.)\|_{L^{\infty}([0,t])}\,.\end{split}$
It concludes the proof of Lemma A.3 by virtue of the assumed regularity of $k$ and $F$, thanks to Lemma A.2 for the first term in $k$. ∎
@@
In the degenerate case where $\nu\equiv 0$, the convergence of $\langle\mu^{N},\varepsilon^{N}\rangle$ to 0 can be directly deduced with the upper-bound of $\varepsilon^{N}$ by $\|k\|_{\infty}\cdot\nu^{N}(\mathcal{Y})$ which converges to 0. In the degenerate case where $\mu\equiv 0$, it suffices to take $3\|k\|_{\infty}\cdot\nu(\mathcal{Y})$ as the uniform upper-bound of $\varepsilon^{N}$ for any $N$ sufficiently large, as $\mu^{N}(\mathcal{X})$ then tends to zero. In the following, we can thus assume that both $\nu(\mathcal{Y})>0$ and $\mu(\mathcal{X})>0$.
The irregularities of $k$ will be located through the following subset $\mathcal{G}[\delta,\eta]$ of $\mathcal{X}\times\mathcal{Y}$, defined for any $\delta,\eta>0$:
@@
while $B[(x,y);\eta]$ denotes the open ball centered in $(x,y)$ of radius $\eta$. The size $\eta$ of the vicinities in (A.3) shall be considered sufficiently small to ensure that discrepancies of order $\delta$ are exceptional. The measurability of the set $\mathcal{G}[\delta,\eta]$ can be more directly verified through the following definition, which happens to be equivalent to the one given in (A.3):
$\mathcal{G}[\delta,\eta]=\cup_{\{q\in{\mathbf{Q}}_{+},m\geq 1\}}k^{-1}([0,q])^{\eta}\cap k^{-1}([q+\delta+2^{-m},\|k\|_{\infty}])^{\eta},$
so that $A^{\eta}$ denotes the $\eta$-vicinity of $A$.
@@
Since $k$ is $(\mu\otimes\nu)$-a.e. continuous, in particular, $k(x,.)$ is $\nu$-a.e. continuous for $x$ on a measurable set $\mathcal{A}\subset\mathcal{X}$ such that $\mu(\mathcal{A})=1$. For any $x\in\mathcal{A}$, thanks to the Portmanteau theorem, see e.g. [29], Subsection IV.3a on the ”Weak Convergence of Probability Measures”, $\varepsilon^{N}(x)$ converges to 0. Remark as compared to the classical version of Portmanteau theorem that we allow $\nu^{N}$ and $\nu$ to be general non-negative finite measure rather than probability measures, given that the proof is not difficult to adapt for this setting. As a consequence of Lebesgue’s dominated convergence theorem, recalling that $\varepsilon^{N}$ is bounded (by 1 under our assumption), we deduce
@@
Step 2: Convergence of $\mu\otimes\nu](\mathcal{G}[\delta,\eta])$ and $\mu\otimes\nu^{N}](\mathcal{G}[\delta,\eta])$ to zero
@@
Note also that the sets $\mathcal{G}[\delta,\eta]$ are increasing as $\delta$ decreases and non-increasing as $\eta$ decreases. Therefore, due to the fact that $k$ is $(\mu\otimes\nu)$-a.e. continuous, the following convergence to zero holds for any $\delta$:
@@
$\mu\otimes\nu^{N}](\mathcal{G}[\delta,\eta])\leq[\mu\otimes\nu](\mathcal{G}[\delta,\eta]^{\eta})+\eta,$ (A.8)
@@
[…]
@@
$\nu\vee\nu^{N}](\mathcal{G}_{x}[\delta,\eta])\geq\delta.$
@@
Step 4: Proof of Proposition 5.9 in the particular case where $(\mu^{N})$, $(\nu^{N})$ are deterministic.
@@
We choose $\eta\in(0,\delta)$ sufficiently small thanks to Step 2, to ensure both that $\mu\otimes\nu](\mathcal{G}[\delta,\eta])$ is smaller than $\delta^{2}/2$ and similarly for $\mu\otimes\nu^{N}](\mathcal{G}[\delta,\eta])$ for any $N$ sufficiently large. Thanks to the Markov inequality, we have
@@
It remains to combine this result with (A.14) and (A.17) to conclude the proof that $(\langle\mu^{N},\varepsilon^{N}\rangle)$ tends to zero as $N$ tends to infinity, since $\delta$ can be taken arbitrarily small. This concludes the proof of Proposition 5.9 in the particular case where the sequences $(\mu^{N})$ and $(\nu^{N})$ are deterministic.
@@
For this final step, we no longer require the sequences $(\mu^{N})$ and $(\nu^{N})$ to be a priori deterministic, though our approach consists in referring to this convenient situation. We exploit [33, Lemma 4.2] to relate the convergence of probability to a.s. convergence of sequence extractions. Let $(N[k])_{k\geq 1}\in\mathbb{N}^{\mathbb{N}}$ be an increasing sequence. Since $\mu^{N}$ and $\nu^{N}$ converge in probability, we can extract a subsequence $(\check{N}[\ell])_{\ell\geq 1}=(N[K[\ell]])_{\ell\geq 1}$ from this sequence such that $(\check{\mu}^{\ell})=(\mu^{\check{N}[\ell]})$ and $(\check{\nu}^{\ell})=(\nu^{\check{N}[\ell]})$ converge a.s. respectively to $\mu$ and $\nu$. On this event of probability 1, we deduce from Step 4 that $\langle\check{\mu}^{\ell},\check{\varepsilon}^{\ell}\rangle$ converges to zero as $\ell$ tends to infinity. Since the convergence of such an extraction of the sequence $(\langle\mu^{N},\varepsilon^{N}\rangle)$ holds whatever the initial extraction, this concludes the proof of Proposition 5.9 in that $(\langle\mu^{N},\varepsilon^{N}\rangle)$ converges in probability to 0. ∎